A parallel beam of light of wavelength 5460 Angstrom is incident at an angle of 30 degree on a plane transmission grating with 6000lines/cm. The highest order of observables spectrum is?
1
Expert's answer
2020-06-08T10:10:15-0400

"d=1/6000=1.67\cdot 10^{-6}m"


"d(\sin\theta_m+\sin\theta_i )=m\lambda"


For "\theta_m=90°"


"m=\frac{d(\sin\theta_m+\sin\theta_i )}{\lambda}=\frac{1.67\cdot10^{-6}(\sin90°+\sin30° )}{5460\cdot10^{-10}}=1.52\approx2"